# Learning Passive Continuous-Time Dynamics with Multistep Port-Hamiltonian Gaussian Processes

### v1

@@ IV-B Vector Field Inference
In this section, we assess how well each method recovers the drift $f(x)$ from noisy, irregularly-sampled trajectories. Fig. 1 shows the performance measured by the mean-squared error (MSE) of the predicted vector field on an evaluation mesh around the simulated true trajectory. After $30$ runs per benchmark and method, we record the MSE and its 95% confidence interval across varying observation noise, time-grid jitter, and Gaussian process ARD initializations. In Fig. 1, we notice that the proposed MS-PHS GP regression with AB-3 integrator performs at least comparable to, if not better than, GP-PHS and MS-ODE in the three benchmarks. Furthermore, we observe that performance typically improves with increasing the order of the integrator.
@@ IV-C Posterior Hamiltonian Recovery
We now move on to inspect the learned Hamiltonian geometry. Fig. 3 overlays the MS-PHS posterior mean and its $\pm 2\sigma_{H}$ band on the ground-truth Duffing surface. Uncertainty expands away from data (black points) and in high-curvature regions of the landscape, while remaining tight near visited states, consistent with the data support and the PHS prior.
Figure 3: 3D Hamiltonian landscape $H$ for the Duffing oscillator. The learned MS-PHS posterior mean $\mu_{H}$ surface (red) is overlaid with a reference/ground-truth surface $H_{\rm true}$ (green) and 96% posterior band $\mu_{H}\pm 2\sigma_{H}$ (translucent blue) for comparison. Black dots mark observed trajectory samples.
@@
Figure 6: Error-uncertainty Ratio of Hamiltonian uncertainty on Duffing. Boxplots show the distribution of the Error-uncertainty ratio $H_{\rm mse}/\sigma^{2}_{H}$ across observation noise levels $\sigma_{x}$, where dashed line represents ideal calibration at 1. MS-PHS-ab-3 remains near or below 1 across $\sigma_{x}$, indicating posterior variance that tracks the error, while GP-PHS-loess-2 drops below 1 as noise increases, evidence of overconfident posterior estimates.
In Fig. 5, GP-PHS-savgol-3 shows error $H_{\mathrm{MSE}}$ increasing rapidly with noise, whereas its reported uncertainty, summarized by $\sigma_{H}^{2}$, lags, especially for $\sigma_{x}^{2}\geq 0.02$. This mismatch is consistent with the decoupling between the LOESS smoothing stage and the GP kernel: the pre-processing can absorb and bias modeling errors that the GP posterior never sees, so the predictive variance fails to reflect the true model error. By contrast, MS-PHS-ab-3 exhibits posterior variance that accurately tracks the realized error. In Fig. 5 (left), $\sigma_{H}^{2}$ rises in step with $H_{\text{MSE}}$ across noise levels, indicating that the posterior dispersion scales with the actual reconstruction difficulty.
To further investigate the issue of calibrating posterior variance to match MSE of Hamiltonian, we leverage the notion of error-uncertainty ratio, defined as $H_{\rm mse}/\sigma^{2}_{H}$. The posterior with $H_{\rm mse}/\sigma^{2}_{H}<1$ is conservative, $H_{\rm mse}/\sigma^{2}_{H}>1$ is overconfident, $H_{\rm mse}/\sigma^{2}_{H}=1$ is well calibrated. The calibration ratio in Fig. 6 makes this explicit: the values concentrate near, or slightly below, the ideal line at $1$ for MS-PHS-ab-3; while GP-PHS-savgol-3 rise above $1$ as noise increases, i.e., uncertainty is underestimated.
TABLE II: Ratio of Hamiltonian posterior variance to MSE ($\sigma_{H}^{2}/H_{\mathrm{MSE}}$) across time-grid jitter levels on Duffing. Entries are “median [Q1, Q3]” across 30 runs.
@@ V Conclusion
We introduced the multistep Port-Hamiltonian Gaussian process (MS-PHS GP) kernel for learning continuous-time dynamics from noisy, irregularly-sampled trajectory data while enforcing port-Hamiltonian structure by construction. By placing a GP prior on the Hamiltonian and embedding variable-step multistep integrator constraints as finite linear functionals, MS-PHS yields closed-form posterior means and variances for both the vector field and the Hamiltonian surface. Theoretically, the construction preserves passivity and compositionality under power-conserving interconnection, and the estimation error decomposes into a multistep truncation term (controlled by integrator order and step size) plus a GP interpolation term (controlled by posterior uncertainty).
Empirically, MS-PHS (particularly with higher-order Adams–Bashforth projectors) recovers vector fields and Hamiltonian geometry more accurately and with better directional alignment than non-physics or prefiltered baselines, while producing well-calibrated posterior uncertainty that tracks empirical error such as $H_{\rm mse}$ across noise and sampling jitter. These properties make MS-PHS especially attractive for downstream control tasks where physically consistent models and reliable uncertainty quantification are essential.

### v2

@@ IV-B Vector Field Inference
In this section, we assess how well each method recovers the drift $f(x)$ from noisy, irregularly-sampled trajectories. Fig. 1 shows the performance measured by the mean-squared error (MSE) of the predicted vector field on an evaluation mesh around the simulated true trajectory. After $30$ runs per benchmark and method, we record the MSE and its 95% confidence interval across varying observation noise, time-grid jitter, and GP ARD initializations. In Fig. 1, we notice that the proposed MS-PHS GP regression with AB-3 integrator performs at least comparable to, if not better than, GP-PHS and MS-ODE in the three benchmarks. Furthermore, we observe that performance typically improves with increasing the order of the integrator.
@@ IV-C Posterior Hamiltonian Recovery
We next inspect the learned Hamiltonian geometry. Fig. 3 overlays the MS-PHS posterior mean and its $\pm 2\sigma_{H}$ band on the ground-truth Duffing surface. Uncertainty expands away from data (black points) and in high-curvature regions of the landscape, while remaining tight near visited states, consistent with the data support and the PHS prior.
Figure 3: 3D Hamiltonian landscape $H$ for the Duffing oscillator. The learned MS-PHS posterior mean $\mu_{H}$ surface (red) is overlaid with a reference/ground-truth surface $H_{\rm true}$ (green) and 95% posterior band $\mu_{H}\pm 2\sigma_{H}$ (translucent blue) for comparison. Black dots mark observed trajectory samples.
@@
Figure 6: Error-uncertainty ratio of Hamiltonian uncertainty on Duffing. Boxplots show the distribution of the $H_{\rm mse}/\sigma^{2}_{H}$ across observation noise levels $\sigma_{x}$, where dashed line represents ideal calibration at 1. MS-PHS-ab-3 remains near or below 1 across $\sigma_{x}$, indicating posterior variance that tracks the error, while GP-PHS-loess-2 rise above 1 as noise increases, evidence of overconfident posterior estimates.
In Fig. 5, GP-PHS-savgol-3 shows error $H_{\mathrm{mse}}$ increasing rapidly with noise, whereas its reported uncertainty, summarized by $\sigma_{H}^{2}$, lags, especially for $\sigma_{x}^{2}\geq 0.02$. This mismatch is consistent with the decoupling between the LOESS smoothing stage and the GP kernel: the pre-processing can absorb and bias modeling errors that the GP posterior never sees, so the predictive variance fails to reflect the true model error. By contrast, MS-PHS-ab-3 exhibits posterior variance that accurately tracks the realized error. In Fig. 5 (left), $\sigma_{H}^{2}$ rises in step with $H_{\rm{mse}}$ across noise levels, indicating that the posterior dispersion scales with the actual reconstruction difficulty.
To further investigate the issue of calibrating posterior variance to match the MSE of Hamiltonian, we leverage the notion of error-uncertainty ratio, defined as $H_{\rm mse}/\sigma^{2}_{H}$. The posterior with $H_{\rm mse}/\sigma^{2}_{H}<1$ is conservative, $H_{\rm mse}/\sigma^{2}_{H}>1$ is overconfident, $H_{\rm mse}/\sigma^{2}_{H}=1$ is well calibrated. The calibration ratio in Fig. 6 makes this explicit: the values concentrate near, or slightly below, the ideal line at $1$ for MS-PHS-ab-3; while GP-PHS-savgol-3 rise above $1$ as noise increases, i.e., uncertainty is underestimated.
TABLE II: Error-uncertainty ratio ($H_{\mathrm{mse}}/\sigma_{H}^{2}$) across time-grid jitter levels on Duffing. Entries are “median [Q1, Q3]” across 30 runs.
@@ V Conclusion
We introduced the MS-PHS GP kernel that enables learning of continuous-time dynamics from noisy, irregular trajectories while enforcing port-Hamiltonian structure. Placing a GP prior on $H$ and embedding variable-step multistep constraints as linear functionals yields closed-form posteriors for both the drift and the Hamiltonian. The construction preserves passivity and composability under power-conserving interconnections and decomposes error into statistical and discretization terms.
Empirically, MS-PHS recovers vector fields and Hamiltonian geometry more accurately and with better directional alignment than non-physics or prefiltered baselines. Meanwhile, MS-PHS produces well-calibrated posterior uncertainty that tracks empirical error such as $H_{\rm mse}$ across noise and sampling jitter. These properties make MS-PHS especially attractive for downstream control tasks where physically consistent models and reliable uncertainty quantification are essential.